During a short interval of time, the acceleration a in m/s2 of an automobile is given by a= Ct+Dt3, where the time t is in seconds. What are the units of C and the dimensions of D.
Answer #1

Given units of acceleration are m/s2

units of time are seconds(s)

From

a=Ct

m/s2 =C*s

C=m/s3

From

a=Dt3

m/s2 =D*s3

D=m/s5